A loud sound made by an animal can be heard over 500 mi away. The power of the sound is 630 watts. Find the number of decibels of sound emitted by the animal. The number of decibels D of a sound can be given by the equation D = 10(log I + 16), where I is the power of the sound measured in watts. Round to the nearest whole number.
Found 2 solutions by Boreal, ikleyn:
Answer by Boreal(15235)

D=10 (log 630+16)
=10(2.8+16)
=188 decibels

Answer by ikleyn(52799)

.
A loud sound made by an animal can be heard over 500 mi away. The power of the sound is 630 watts.
Find the number of decibels of sound emitted by the animal. The number of decibels D of a sound
can be given by the equation D = 10(log I + 16), where I is the power of the sound measured in watts.
Round to the nearest whole number.
~~~~~~~~~~~~~~~


In your post, double check this absurdist value of 500 miles away, where an animal sound can be heard.

I think, it should be 500 meters, instead of 500 miles.